# Enhancing Answer Reliability Through Inter-Model Consensus of Large Language Models

### v1

@@ 4 Results
4.1 Inter-Model Consistency Analysis
We assessed the level of agreement among the models—GPT-4, LLaMA, Gemini, and Claude—for each generated question. In each experiment, one model generated $100$ questions, and the remaining three independently provided answers without collaboration. The implantation and data are available here Amiri-Margavi (2024).
@@ 4.1 Inter-Model Consistency Analysis
[…]
These findings suggest that inter-model consistency depends significantly on the question-generating model. The higher agreement levels observed with Claude and GPT-4 may be attributed to their advanced training and more sophisticated understanding of complex statistical concepts. Conversely, the lower consensus rates with LLaMA and Gemini could be influenced by differences in their model architectures, training data, or interpretation nuances. This variation underscores the impact of the question generator on the collaborative performance of LLMs in specialized domains.
To evaluate the reliability and precision of answers generated by LLM) in scenarios lacking ground-truth data, we employed two complementary mechanisms: majority voting and confidence interval analysis. These approaches quantify inter-model agreement and assess alignment with a reference answer implicitly provided by the question-generating LLM.
Using the majority voting approach, an answer is considered reliable if at least two other LLMs agree with the question-generating LLM’s response. Figure 2 presents the majority vote and reliability percentages for various LLMs. Key observations reveal that majority vote percentages are consistently higher than reliability percentages across all models, indicating strong inter-model agreement but fewer instances where the models’ answers align with the reference answer from the question-generating LLM.
@@
[…]
This analysis highlights the utility of majority voting in quantifying inter-model agreement and reliability. The gap between majority vote and reliability percentages highlights the challenges of aligning responses with a reference answer, particularly in the absence of ground truth data. These findings demonstrate the effectiveness of majority voting as a plausible mechanism for validating answers and assessing model performance when definitive correctness cannot be directly measured.
To further evaluate the precision and reliability of inter-model agreement, we calculated confidence intervals (CIs) for the consensus rates of each model. Confidence intervals provide a statistical range within which the consensus rate is likely to fall, enabling an assessment of variability and robustness. Table 3 summarizes each model’s lower and upper bounds of the CIs.
@@
Analysis of the confidence intervals reveals distinct patterns in the precision and reliability of the models. Claude demonstrates the narrowest CI (0.80–0.93), suggesting high reliability and minimal variability in its consensus rates. GPT-4 follows closely with a CI of (0.75–0.90), indicating similarly high reliability but slightly increased variability. In contrast, Gemini and LLaMA exhibit broader CIs, spanning (0.60–0.78) and (0.55–0.74), respectively, reflecting more significant variability and reduced precision in their consensus rates.
These findings provide quantitative evidence of inter-model performance differences. Models with narrower CIs, such as Claude and GPT-4, offer more reliable consensus rates, making them more suitable for high-precision tasks. Conversely, the wider CIs observed for Gemini and LLaMA highlight potential inconsistencies and need further refinement. Confidence interval analysis thus offers a robust framework for assessing model agreement and reliability without ground truth data.
4.2 Statistical Significance Testing
@@
The chi-square test was performed to determine whether the observed agreement among the models occurred by chance. The null hypothesis assumed that the distribution of answers was random, indicating no meaningful consensus. The test statistic was calculated based on eq: 6, in which the number of choices $K=4$ and $n_{j}=3$ is the number of answering models. The p-values obtained from the test are summarized in Table 4.
@@
All p-values were below the standard significance level ($\alpha=0.01$), leading us to reject the null hypothesis. The extremely small p-values for Claude, Gemini, and LLaMA indicate statistically solid significance, confirming that their agreements are highly unlikely to be due to random chance. GPT-4 also shows statistical significance, though its p-value is less extreme than the others. These results support the hypothesis that the models’ agreements reflect meaningful consensus rather than random behavior.
Fleiss’Kappaa was further employed to evaluate inter-rater agreement among the models. This metric provides a robust measure of consistency in responses. Table 5 presents the kappa values and their interpretations.
@@
The kappa values reveal varying levels of agreement across the models. Claude demonstrated substantial agreement, reflecting strong consistency with other models. GPT-4 and LLaMA Combiningderate agreement, indicating reasonable consistency but some response variability. Conversely, Gemini showed fair agreement, suggesting noticeable differences in its responses compared to the other models.
@@
These results highlight potential differences in model architectures, training data, or reasoning approaches. For instance, Gemini’s lower agreement could stem from less accurate or diverse data sources, while Claude’s substantial agreement may be attributed to robust training on well-aligned datasets. By combining chi-square testing and Fleiss’ Kappa analysis, we comprehensively evaluate the statistical significance and inter-model agreement, demonstrating the reliability and coherence of LLM-generated answers.
5 Discussion
5.1 Enhancing Answer Reliability Through Collaboration
The high consensus rate among the models, as demonstrated in both the majority vote and reliability analyses, underscores the potential of collaborative dynamics in enhancing answer reliability. The substantial Fleiss’ Kappa values confirm that the agreement among the models is statistically significant and not due to chance. This collective agreement suggests that the models possess a shared understanding of statistical concepts, even without ground-truth answers.
In scenarios where expert validation is unavailable—such as rapidly evolving domains or complex problem-solving contexts—leveraging multiple LLMs for consensus-based validation provides a valuable proxy for correctness. For instance, questions where Claude and GPT-4 exhibit high agreement can be considered more reliable, whereas questions with lower agreement or broader confidence intervals (e.g., those involving LLaMA) may require further assessment. The p-values from the chi-square test further support this analysis, allowing researchers to identify questions where models show strong alignment versus those with more significant variability. This insight can guide future efforts to refine questions with high disagreement rates, which may be inherently ambiguous or beyond the current capabilities of LLMs.
5.2 Implications for AI and Education
Our findings significantly impact applying LLMs in AI-driven education and content generation. In collaborative learning, LLMs can enhance the quality of educational content by validating and refining answers collaboratively, even without human-provided ground-truth answers. This capability could revolutionize automated learning systems, enabling them to independently generate and validate high-quality content. Such systems could provide immediate, reliable feedback to learners, adapt to new information swiftly, and support educators by handling routine assessment tasks.
5.3 Limitations
This study acknowledges several limitations that may impact the findings and interpretations. One significant limitation is the potential ”model similarities” or ”model homogeneity.” Despite differences in architecture and training, the analyzed LLMs may share overlapping training data, leading to correlated errors and overestimating consensus. This overlap could result in models reinforcing each other’s misconceptions rather than providing independent validation.
A second limitation is the ”lack of a human benchmark.” Without expert human evaluation, it remains challenging to assess the absolute correctness of the consensus answers definitively. The reliability metrics employed, while meaningful, serve as subjective proxies and cannot fully substitute for expert judgment.
Third, the study highlights the risk of ”bias propagation.” If all models share similar misconceptions, they may propagate and reinforce incorrect answers through agreement, which could be particularly problematic in specialized or sensitive domains where accuracy is critical.
Finally, this study is constrained by its ”static evaluation” nature. It captures a snapshot of the models’ capabilities at a single point in time and does not account for potential updates, improvements, or degradations in model performance over time. Future work should consider dynamic evaluation to track evolving model capabilities better.
5.4 Conclusion
[…]
These findings prove that collaborative interactions among large language models (LLMs) can improve response reliability through consensus mechanisms. The study highlights the potential of these methodologies for applications in AI-driven education, automated validation systems, and collaborative AI frameworks. For example, leveraging highly reliable models such as Claude and GPT-4 in cooperative settings could enhance the accuracy and trustworthiness of AI-generated responses in high-stakes domains.
Future research should explore integrating models with diverse architectures and training datasets to mitigate the risk of correlated errors and investigate the impact of model heterogeneity on consensus rates. Further analysis of bias propagation is also critical to ensure collaborative systems do not inadvertently reinforce shared misconceptions among models. Addressing these directions will significantly improve consensus-based validation frameworks’ robustness, scalability, and reliability, laying the groundwork for more effective and trustworthy AI systems.

### v2

@@ 4 Results
We evaluated the level of agreement among the models—GPT-4, LLaMA, Gemini, and Claude—for each generated question. In each experiment, one model generated $100$ questions, and the remaining three independently provided answers without collaboration. The data are available here Amiri-Margavi (2024).
Inter-Model Consistency Analysis
@@
Figure 1 illustrates the levels of agreement between the models based on their responses to the questions generated by different LLMs. The results indicate that Claude and GPT-4 achieved the highest full agreement rates, with $86\%$ and $82\%$ of the responses aligning when they served as the question generator. In contrast, Gemini and LLaMA showed lower consensus, with full agreement rates of $70\%$ and $65\%$, respectively. This suggests that the questions generated by Claude are more comprehensible to other LLMs, as evidenced by the consistent selection of the same answer among multiple choice options $\{A,B,C,D\}$. In contrast, when LLaMA generated the questions, only $65\%$ of the responses reached full agreement, indicating greater variability in interpretation.
Further analysis supports these observations. LLaMA-generated questions resulted in a partial agreement rate of $27\%$ and the highest non-agreement rate of $8\%$, suggesting that their questions may be more ambiguous or challenging to interpret. In contrast, Claude’s questions had a partial agreement rate of only $14\%$ and no cases of complete disagreement, supporting the hypothesis that their ability to generate well-structured and comprehensible questions was compared to the other LLMs, particularly LLaMA.
To evaluate the reliability and precision of the answers generated by LLM in scenarios lacking ground-truth data, we used two complementary mechanisms: majority voting and confidence interval analysis. These approaches quantify inter-model agreement and assess alignment with a reference answer provided by the question-generating LLM.
Using the majority voting approach, an answer is considered reliable if at least two other LLMs agree with the question-generating LLM’s response. Figure 2 presents the percentages of majority vote and reliability of various LLMs. The results indicate that the percentages of majority votes remain high in all models, with Claude achieving the highest at $100\%$, followed by Gemini and GPT-4 at $98\%$, and LLaMA at $92\%$. This suggests that in most cases, two or three LLMs consistently selected the same answer among multiple choice options $\{A,B,C,D\}$.
Reliability analysis reveals notable differences in the frequency with which models align with the reference answer. Claude demonstrates the highest reliability percentage ($92\%$), followed closely by GPT-4 ($90\%$) and Gemini ($88\%$), suggesting strong consistency with the LLM response generating the questions. In contrast, LLaMA exhibits the lowest reliability percentage ($77\%$), indicating a reduced alignment with reference answers.
@@
To further evaluate the precision and reliability of inter-model agreement, we calculated confidence intervals (CIs) for the consensus rates of each model. Confidence intervals provide a statistical range within which the consensus rate is likely to fall, allowing an assessment of variability and robustness. Table 3 summarizes the lower and upper bounds of each model of the CIs.
@@
Analysis of the confidence intervals reveals distinct patterns in the precision and reliability of the models. Claude demonstrates the narrowest CI (0.800.800.800.800.930.930.930.93), indicating high reliability and minimal variability in its consensus rates. GPT-4 closely follows a CI of (0.750.750.750.750.900.900.900.90), suggesting similarly strong reliability but with slightly greater variability. In contrast, Gemini and LLaMA exhibit broader confidence intervals (0.600.600.600.600.780.780.780.78) and (0.550.550.550.550.740.740.740.74), respectively, reflecting higher variability and reduced precision in their consensus rates.
Furthermore, the low upper bounds of LLaMA ($0.74$) and Gemini ($0.78$) indicate that even in the most favorable scenarios, these models achieve only moderate agreement levels with other LLMs. In contrast, Claude ($0.80$) and GPT-4 ($0.75$) maintain high lower bounds, which means that even in the least favorable cases, they still uphold strong agreement among models. This consistency suggests that Claude and GPT-4 generate questions that are better structured and more interpretable, allowing for greater inter-model alignment. However, Gemini and LLaMA show the need for further optimization to improve inter-model agreement and reliability in question generation.
Statistical Significance Testing
@@
The chi-square test was performed to determine whether the observed agreement among the models occurred by randomness. The null hypothesis assumed that the distribution of the responses was random, indicating that there was no meaningful consensus. The test statistic was calculated based on eq: 6, in which the number of choices $K=4$ and $n_{j}=3$ is the number of answering models. The p-values obtained from the test are summarized in Table 4.
@@
All $p$-values were below the standard significance level ($\alpha=0.01$), which led us to reject the null hypothesis. The extremely small $p$-values for Claude, Gemini, and LLaMA indicate statistically solid significance, confirming that their agreements are highly unlikely to be due to randomness. GPT-4 also shows statistical significance, although its $p$-value is less extreme than the others. These results support the hypothesis that the models’ agreements reflect meaningful consensus rather than random behavior.
Fleiss’ Kappa was further employed to evaluate inter-rater agreement among the models. This metric provides a robust measure of consistency in responses. Table 5 presents the kappa values and their interpretations.
@@
The kappa values reveal varying levels of agreement across the models. Claude demonstrated substantial agreement, reflecting strong consistency with other models. GPT-4 and LLaMA have a moderate agreement, indicating reasonable consistency but some variability in response. Conversely, Gemini showed fair agreement, suggesting noticeable differences in its responses compared to the other models.
@@
These results highlight potential differences in model architectures, training data, fine-tuning processes, or reasoning approaches. For instance, Gemini’s lower agreement could stem from less accurate or diverse data sources, while Claude’s substantial agreement may be attributed to robust training on well-aligned datasets. By combining chi-square testing and Fleiss’s Kappa analysis, we comprehensively evaluate the statistical significance and inter-model agreement, demonstrating the reliability and coherence of LLM-generated answers.
Conclusion
This study demonstrates that collaborative dynamics among multiple LLMs can significantly enhance response reliability even in the absence of ground-truth data. We quantified the agreement between models using consensus-based validation and confidence interval analysis, revealing distinct strengths and weaknesses across models. Our analysis emphasizes the crucial role of the question-generating model in shaping inter-model agreement: Claude consistently produced well-structured, comprehensible questions, while LLaMA’s output exhibited greater ambiguity. In addition, Claude and GPT-4 emerged as the most reliable collaborators, achieving reliability percentages of $92\%$ and $90\%$, respectively, with narrow confidence intervals (Claude: 0.800.800.800.800.930.930.930.93, GPT-4:$0.75$-$0.90$), indicating high precision and consistency. In contrast, Gemini showed intermediate performance with an $88\%$ reliability rate and a confidence interval of $0.60$-$0.78$, whereas LLaMA recorded the lowest reliability at $77\%$ with the widest confidence interval (0.550.550.550.550.740.740.740.74), reflecting significant variability. The statistical significance of these results, as evidenced by a very small $p$-value, confirms that the observed inter-model agreement is not random, thus reinforcing the robustness of the consensus mechanism in identifying trustworthy responses.
Future research should integrate human experts into the validation process to establish a reliable ground truth for evaluating majority vote approaches. Our findings emphasize that LLMs are not equally effective in all tasks, indicating that improvements in question-generation strategies and fine-tuning are necessary to improve consistency and reduce ambiguity. This study acknowledges several limitations, notably the potential for model homogeneity. Despite differences in architecture and training, analyzed LLMs can share overlapping training data, leading to correlated errors and an overestimation of consensus, as models can reinforce misinterpretation of each other rather than provide independent validation. In addition, further investigation into bias propagation is critical to ensure that collaborative systems do not inadvertently perpetuate shared misconceptions. Addressing these challenges will enhance the robustness, scalability, and reliability of consensus-based validation frameworks, ultimately laying the foundation for more effective and trustworthy AI systems.